# Training-Conditional Coverage Bounds under Covariate Shift

### v1

@@ Abstract
Training-conditional coverage guarantees in conformal prediction concern the concentration of the error distribution, conditional on the training data, below some nominal level. The conformal prediction methodology has recently been generalized to the covariate shift setting, namely, the covariate distribution changes between the training and test data. In this paper, we study the training-conditional coverage properties of a range of conformal prediction methods under covariate shift via a weighted version of the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality tailored for distribution change. The result for the split conformal method is almost assumption-free, while the results for the full conformal and jackknife+ methods rely on strong assumptions including the uniform stability of the training algorithm.
@@ 1 Introduction
Conformal prediction is a framework for constructing distribution-free predictive confidence regions as long as the training and test data are exchangeable Vovk et al. (2005) (also see Shafer and Vovk (2008); Vovk et al. (2009); Vovk (2012)). Specifically, let $((X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),(X_{\text{test}},Y_{\text{test}}))$ denote a tuple of exchangeable data points, consisting of a training sequence of $n$ samples ${\bf D}_{n}:=((X_{1},Y_{1}),\ldots,(X_{n},Y_{n}))\in(\mathcal{X}\times\mathcal{Y})^{n}$ and one test sample $(X_{\text{test}},Y_{\text{test}})\in\mathcal{X}\times\mathcal{Y}$. The conformal prediction provides a coverage of $Y_{\text{test}}$ in the sense of
@@
where $\hat{C}_{\alpha}:\mathcal{X}\to 2^{\mathcal{Y}}$ is a data-dependent map. This type of guarantee is referred to as marginal coverage, as it is averaged over all the training and test data.
One natural direction to stronger results is to devise conditional coverage guarantee
@@
However, it has been shown in Vovk (2012); Foygel Barber et al. (2021); Lei and Wasserman (2014) that it is impossible to obtain (non-trivial) distribution-free prediction regions $\hat{\mathcal{C}}(x)$ in the finite-sample regime; relaxed versions of this type of guarantee have been extensively studied (see Jung et al. (2022); Gibbs et al. (2023); Vovk et al. (2003) and references therein). As an alternative approach, several results (e.g., Vovk (2012); Bian and Barber (2023)) have been reported on the training-conditional guarantee by conditioning on ${\bf D}_{n}$, which is also more appealing than the marginal guarantee as can be seen below. Define the following miscoverage rate as a function of the training data,
$P_{e}({\bf D}_{n}):=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}(X_{\text{test}})|{\bf D}_{n}).$
Note that the marginal coverage in (1) is equivalent to $\operatorname{\mathbb{E}}[P_{e}({\bf D}_{n})]\leq\alpha$. The training-conditional guarantees are of the following form, for some small $\delta>0$,
@@
For the $K$-fold CV+ with $m$ samples in each fold, the conditional coverage bound
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})\geq 2\alpha+\sqrt{2\log(K/\delta)/m}\right)\leq\delta$ (2)
@@
is established in Bian and Barber (2023). They have also shown that distribution-free training-conditional guarantees for full conformal and jackknife+ methods are impossible without further assumptions; in particular, they conjectured that a certain form of algorithmic stability is needed for full conformal and jackknife+. Recently, Liang and Barber (2023) proposed conditional coverage bounds for jackknife+ and full conformal prediction sets under the assumption that the training algorithm is symmetric and satisfies certain stability conditions (see Section 4 for more details). In this line of research, samples are assumed to be i.i.d., which is not only exchangeable but also ergodic and admits some nice concentration properties. However, this assumption can be violated in the application. In particular, the input data distribution during deployment can differ from the distribution observed during training. This phenomenon is called distribution shift and it is a crucial problem in trustworthy machine learning (see Section 2.3). In this regard, split and full conformal prediction methods as the central methods for distribution-free uncertainty quantification of black-box models have recently been extended to handle a popular type of distribution shift called covariate shift Tibshirani et al. (2019). In the covariate shift setting, the distribution of covariates in the test data differs from the one observed in the training data, but, the conditional distribution of the response given the features remains the same across the training and test populations. A similar extension has been made for the jackknife+ method Prinster et al. (2022).
@@
In this paper, we study the training-conditional coverage properties of the conformal prediction methods under covariate shift. We present concentration bounds of the training-conditional errors for weighted jackknife+ (JAW) Prinster et al. (2022), full, and split conformal methods. In terms of distributions, we assume the existence of $dQ/dP$ and consider two regimes: (1) $\|dQ/dP\|_{\infty}<\infty$ and (2) $\|dQ/dP\|_{P,2}<\infty$, where $Q$ and $P$ denote the distribution of the test and training data, respectively. These two settings have also been considered in Ma et al. (2023) to study the optimality of the reproducing kernel Hilbert space (RKHS) regressions under the covariate shift setting, where it is assumed that the true regression function (i.e., $\operatorname{\mathbb{E}}(Y|X=x)$) belongs to the RKHS. Regarding the training algorithm, no assumption is made for the split conformal method. However, full conformal and jackknife+ methods are analyzed under the assumption of uniform stability as explained below. Let $\hat{\mu}_{{\bf D}_{n}}$ denote the predictor function estimated using the training data ${\bf D}_{n}$. A training algorithm is called uniformly stable if,
@@ 2.1 Full conformal and split conformal
where $\hat{F}^{-1}_{S}(1-\alpha)$ denotes the empirical quantile function with respect to $S$. Thus,
@@ 2.2 Jackknife+
where $\hat{\mu}=T({\bf D}_{n})$ and $\mathcal{S}^{\text{J}}:=\{s(X_{i},Y_{i};\hat{\mu}^{-i}):1\leq i\leq n\}$ and $\hat{\mu}^{-i}:=T(((X_{1},Y_{1}),\ldots,(X_{i-1},Y_{i-1}),(X_{i+1},Y_{i+1}),\ldots,(X_{n},Y_{n})))$. Despite its effectiveness, no general finite-sample guarantees are known for jackknife. Recently, Barber et al. (2021) proposed jackknife+, a modified version of the jackknife for $\mathcal{Y}=\mathbb{R}$ and $f(y,y^{\prime})=|y-y^{\prime}|$, and established $(1-2\alpha)$ coverage lower bound for it. Let
@@
Also, the jackknife+ has been generalized to CV+ for $K$-fold cross-validation, and $(1-2\alpha-\sqrt{2/|\mathcal{D}^{\text{train}}|})$ coverage lower bound is established.
@@ 2.3 Conformal prediction under distribution shift
To handle the covariate shift, a weighted version of conformal prediction was first proposed in the seminal work by Tibshirani et al. Tibshirani et al. (2019). The key assumption is that the ratio of test to training covariate, $dQ/dP$, needs to be known; we follow the same assumption in this work. This reweighting scheme has been extended to various settings in conformal prediction Podkopaev and Ramdas (2021); Lei and Candès (2021); Fannjianga et al. ; Guan (2023). The general domain adaption problem has also been an active area from a causal perspective (e.g.,  Zhang et al. (2013a); Peters et al. (2016); Gong et al. (2016); Chen and Bühlmann (2021); Du and Xiang (2023)).
@@
$F_{Q,n}(t)=w_{\text{test}}\boldsymbol{1}\{t=\infty\}+\sum_{i\in\mathcal{I}}w_{i}\boldsymbol{1}\{s(Z_{i})\leq t\}.$
@@ 3.1 Split conformal
$\operatorname{\mathbb{P}}\left(P_{e}^{\text{split}}({\bf D}_{n})>\alpha+\bigl{(}\sqrt{2B\log(4/\delta)}+3C\bigr{)}\sqrt{\frac{B}{m}}\right)\leq\delta.$
@@
Remark 1
Relaxing the assumption $dQ/dP\leq B$ to $\|dQ/dP\|_{P,2}\leq B$, we get
$\operatorname{\mathbb{P}}\left(P_{e}^{\text{split}}({\bf D}_{n})>\alpha+\frac{2B(3C+1)}{\delta\sqrt{m}}\right)\leq\delta,\qquad\ 0<\delta\leq 1.$ (8)
For simulations, we refer the readers to the experiments on Airfoil data Tibshirani et al. (2019), which shows the empirical distribution of $P_{e}^{\text{split}}({\bf D}_{n})$. Note that the theoretical guarantees in Tibshirani et al. (2019) only concern the marginal coverage, e.g., $\operatorname{\mathbb{E}}[P_{e}^{\text{split}}({\bf D}_{n})]=\operatorname{\mathbb{P}}(Y_{\text{test}}\not\in\hat{C}^{\text{split}}_{\alpha}(X_{\text{test}}))$ for split conformal.
@@ 3.2 Full conformal and jackknife+
Let $\mu_{\beta}:\mathcal{X}\to\mathbb{R}$ denote a predictor function parameterized by $\beta\in\mathbb{R}^{p}$. By a slight abuse of notation, let the map $T:\cup_{n\geq 1}(\mathcal{X}\times\mathcal{Y})^{n}\to\mathbb{R}^{p}$ denote a training algorithm for estimating $\beta$, hence, $\hat{\beta}_{n}=T({\bf D}_{n})$ In this case, we have $\hat{\mu}_{{\bf D}_{n}}=\mu_{\hat{\beta}_{n}}$.
@@
$\operatorname{\mathbb{P}}\Bigg{(}P_{e}^{\,\text{J+}}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log{4}/{\delta}}+3C\right)\sqrt{\frac{B}{n}}+L_{Q,n-1}\,A(n,p,\epsilon)\Bigg{)}\leq\epsilon+\delta,$
@@
Using the same arguments as in the proof of this theorem, one can get a coverage bound for the CV+ as well. Unlike (2) which is meaningful only if the number of samples in each fold $m$ is large, the bound we present in the following corollary is suitable for cases where $m/n\to 0$.
@@
$\operatorname{\mathbb{P}}\Bigg{(}$ $P_{e}^{\,\text{CV+}}({\bf D}_{n})>\alpha+\sqrt{\frac{\log(2/\delta)}{2n}}+2\,m\,L_{n-m}\,\kappa_{2}\,c_{n-m}\bigg{(}\frac{1}{\kappa_{1}}+\sqrt{\frac{n}{2\kappa_{1}^{2}}\log\frac{2p}{\epsilon}}\bigg{)}\Bigg{)}\leq\epsilon+\delta.$
@@
$\operatorname{\mathbb{P}}\left(P_{e}^{\text{full}}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log(4/\delta)}+3C\right)\sqrt{\frac{B}{n}}+L_{Q,n}\,E(n,p,\epsilon)\right)\leq\epsilon+\delta.$
@@ 4 Discussion
where $\hat{\mu}_{{\bf D}_{n}}=T({\bf D}_{n})$ and $X_{\text{test}}\perp\!\!\!\perp{\bf D}_{n+m}$ with ${\bf D}_{n+m}=((X_{1},Y_{1}),...,(X_{n+m},Y_{n+m}))$. Although weaker than uniform stability, these parameters are yet not well-understood in a practical sense. To elaborate on the difference between this approach and uniform stability, we evaluate these bounds for the ridge regression under exchangeability.
@@
$\operatorname{\mathbb{P}}\Bigg{(}P_{e}^{\text{full}}({\bf D}_{n})>\alpha+O\big{(}n^{-1/2}\big{(}\sqrt{{\log(1/\delta)}}+$ $\sqrt{p\log(2p/\epsilon)}\big{)}\big{)}\Bigg{)}\leq\epsilon+\delta.$

### v2

@@ Abstract
Conformal prediction methodology has recently been extended to the covariate shift setting, where the distribution of covariates differs between training and test data. While existing results ensure that the prediction sets from these methods achieve marginal coverage above a nominal level, their coverage rate conditional on the training dataset—referred to as training-conditional coverage—remains unexplored. In this paper, we address this gap by deriving upper bounds on the tail of the training-conditional coverage distribution, offering probably approximately correct (PAC) guarantees for these methods. Our results quantify the relationship between the quality of the prediction sets and the severity of distributional changes, and can potentially be used to compute more efficient prediction sets.
@@ 1 Introduction
Conformal prediction is a framework for constructing distribution-free and model-agnostic predictive confidence regions under the exchangeability assumption for the training and test samples Vovk et al. (2005) (also see Shafer and Vovk (2008); Vovk et al. (2009); Vovk (2012)). Let $((X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n}),(X_{\text{test}},Y_{\text{test}}))$ denote a tuple of exchangeable data points, consisting of a training sequence of $n$ samples ${\bf D}_{n}:=((X_{1},Y_{1}),\ldots,(X_{n},Y_{n}))\in(\mathcal{X}\times\mathcal{Y})^{n}$ and one test sample $(X_{\text{test}},Y_{\text{test}})\in\mathcal{X}\times\mathcal{Y}$. For a fixed error rate level $\alpha$, the conformal prediction framework, provides a prediction set $\hat{C}_{\alpha}(X_{\text{test}})$ for $Y_{\text{test}}$ that satisfies
@@
where $\hat{C}_{\alpha}:\mathcal{X}\to 2^{\mathcal{Y}}$ is a data-dependent map. This type of guarantee is referred to as marginal coverage, as it is averaged over both the training and test data.
One natural direction to stronger results is to provide a coverage guarantee conditional on the test data point $X_{\text{test}}$, i.e.,
@@
However, when $X_{\text{test}}$ has a continuous distribution, it has been shown in Vovk (2012); Foygel Barber et al. (2021); Lei and Wasserman (2014) that it is impossible to obtain (non-trivial) distribution-free prediction regions $\hat{\mathcal{C}}_{\alpha}(x)$ in the finite-sample regime; relaxed versions of this type of guarantee have been extensively studied (see Jung et al. (2022); Gibbs et al. (2023); Vovk et al. (2003) and references therein). As a different approach to stronger guarantees, several results (e.g., Vovk (2012); Bian and Barber (2023)) have been reported on the training-conditional guarantee by conditioning on ${\bf D}_{n}$, which is also more appealing than the marginal guarantee as can be seen below. Define the following miscoverage rate as a function of the training data,
$P_{e}({\bf D}_{n}):=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}(X_{\text{test}})|{\bf D}_{n}).$
Note that the marginal coverage in (1) is equivalent to $\operatorname{\mathbb{E}}[P_{e}({\bf D}_{n})]\leq\alpha$. The training-conditional guarantees concern the concentration of the conditional error rate below the nominal level $\alpha$ and they have the following form, for some small $\delta>0$,
@@
Recently, Barber et al. (2021) proposed modified versions of jackknife and cross validation (CV), namely jackknife+ and CV+, which can be used to compute conformal prediction sets. For the $K$-fold CV+ with $\ell$ samples in each fold, the training-conditional coverage bound
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})\geq 2\alpha+\sqrt{2\log(K/\delta)/\ell}\right)\leq\delta$ (2)
@@
is established in Bian and Barber (2023). They have also shown that distribution-free training-conditional guarantees for full conformal and jackknife+ methods are impossible without further assumptions; in particular, they conjectured that a certain form of algorithmic stability is needed for full conformal and jackknife+. Recently, Liang and Barber (2023) proposed conditional coverage bounds for jackknife+ and full conformal prediction sets under the assumption that the training algorithm is symmetric and satisfies certain stability conditions (see Section 5 for more details). In this line of research, samples are assumed to be i.i.d., which is not only exchangeable but also ergodic and admits some nice concentration properties. However, this assumption can be violated in the application. In particular, the input data distribution during deployment can differ from the distribution observed during training. This phenomenon is called distribution shift and it is a crucial problem in trustworthy machine learning (see Section 2.3). In this regard, split and full conformal prediction methods as the central methods for distribution-free uncertainty quantification of black-box models have recently been extended to handle a popular type of distribution shift called covariate shift Tibshirani et al. (2019). In the covariate shift setting, the distribution of covariates in the test data differs from the one observed in the training data, but, the conditional distribution of the response given the features remains the same across the training and test populations. A similar extension has been made for the jackknife+ method Prinster et al. (2022). However, these results primarily focus on marginal coverage. Although the weighted conformal prediction is guaranteed to keep the marginal coverage rate above the nominal level, it often reduces the concentration of the training-conditional coverage rate $\operatorname{\mathbb{P}}(Y_{\text{test}}\in\hat{C}_{\alpha}(X_{\text{test}})|{\bf D}_{n})$ (i.e., coverage rate conditional on the training data ${\bf D}_{n}$) around the nominal level $1-\alpha$. In particular, the weighting scheme leads to heavier tails for the training-conditional coverage rate. This phenomenon is illustrated in Figure 1, where ordinary and weighted conformal methods with $80\%$ target coverage rate ($\alpha=0.2$) are used for constructing prediction intervals under exchangeability and distribution shift settings, respectively. The distribution shift is introduced artificially via resampling according to an exponential tilt as in Section 2.3 from Tibshirani et al. (2019). It can be observed that the tails of the training conditional coverage get heavier when the weighting scheme is used to handle the distribution shift. In this paper, we explore the quality of the weighted conformal prediction regions by computing upper bounds on the tails of the training-conditional coverage distribution, quantifying the relationship between the tail behavior and the extent of distribution change. In other words, we examine the efficiency of weighted conformal prediction methods for a generic training dataset under distribution shift.
Figure 1: Histograms of the training-conditional coverage rates are presented for four datasets from the UCI Machine Learning Repository: Wine Quality (top left), Abalone (top right), Concrete Compressive Strength (bottom left), and Combined Cycle Power Plant (bottom right). Models were trained using 1000 samples and an RKHS regression model with a Gaussian kernel.
@@
We present concentration bounds of the training-conditional errors for weighted jackknife+ (JAW) Prinster et al. (2022), full, and split conformal methods. Regarding the training algorithm, no assumption is made for the split conformal method. However, full conformal and jackknife+ methods are analyzed under the assumption of uniform stability as explained below. Let $\hat{\mu}_{{\bf D}_{n}}$ denote the predictor function estimated using the training data ${\bf D}_{n}$. A training algorithm is called uniformly stable if,
@@ 2.1 Full conformal and split conformal
where $\hat{F}^{-1}_{S}(\cdot)$ denotes the empirical quantile function with respect to $S$. Thus,
@@ 2.2 Jackknife+
where $\hat{\mu}=T({\bf D}_{n})$, $\mathcal{S}^{\text{J}}:=\{s(X_{i},Y_{i};\hat{\mu}^{-i}):1\leq i\leq n\}$ and $\hat{\mu}^{-i}:=T(((X_{1},Y_{1}),\ldots,(X_{i-1},Y_{i-1}),(X_{i+1},Y_{i+1}),\ldots,(X_{n},Y_{n})))$. Despite its effectiveness, no general finite-sample guarantees are known for jackknife. Recently, Barber et al. (2021) proposed jackknife+, a modified version of the jackknife for $\mathcal{Y}=\mathbb{R}$ and $f(y,y^{\prime})=|y-y^{\prime}|$, and established $(1-2\alpha)$ finite-sample coverage lower bound for it. Let
@@
Also, the jackknife+ has been generalized to CV+ for $K$-fold cross-validation, and $(1-2\alpha-\sqrt{2/n})$ coverage lower bound is established.
@@ 2.3 Conformal prediction under distribution shift
To handle the covariate shift, a weighted version of conformal prediction was first proposed in the seminal work by Tibshirani et al. (2019). The key assumption is that the likelihood ratio $dQ/dP=dQ_{X}/dP_{X}$, needs to be known; we follow the same assumption in this work. This reweighting scheme has been extended to various settings in conformal prediction Podkopaev and Ramdas (2021); Lei and Candès (2021); Fannjianga et al. ; Guan (2023). The general domain adaption problem has also been an active area from a causal perspective (e.g.,  Zhang et al. (2013a); Peters et al. (2016); Gong et al. (2016); Chen and Bühlmann (2021); Du and Xiang (2023)).
@@
$F_{Q,|\mathcal{I}|}(t)=w_{\text{test}}\boldsymbol{1}\{t=\infty\}+\sum_{i\in\mathcal{I}}w_{i}\boldsymbol{1}\{s(Z_{i})\leq t\}.$
@@ 3.1 Split conformal
$\operatorname{\mathbb{P}}\left(P_{e}^{\text{split}}({\bf D}_{n})>\alpha+\bigl{(}\sqrt{2B\log(4/\delta)}+2C\bigr{)}\sqrt{\frac{B}{m}}\right)\leq\delta.$
@@
Corollary 1
Relaxing the assumption $dQ/dP\leq B$ to $\|dQ/dP\|_{P,2}\leq B$ via Lemma 2, we get
$\operatorname{\mathbb{P}}\left(P_{e}^{\text{split}}({\bf D}_{n})>\alpha+\frac{2B(2C+1)}{\delta\sqrt{m}}\right)\leq\delta,\qquad\ 0<\delta\leq 1.$ (8)
To assess the robustness of weighted conformal prediction methods, we compute training-conditional error bounds with no assumptions on $dQ/dP$. The proof follows from Lemma 4 and the proof of Theorem 1.
Theorem 2
Fix $B,\delta>0$ and $\epsilon\in(0,1/2)$. It holds that,
$P_{e}^{\text{split}}({\bf D}_{n})<\alpha+3\delta+4\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>m\right)+C^{\prime}\sqrt{\frac{B}{m^{\prime}}}+\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right),$
with probability at least,
$1-4e^{-m^{\prime}\delta^{2}/(2B^{2})}$ $-2e^{-2m^{2\epsilon}}-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>m\right)$
$-\frac{8}{m\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq m\right\}\right)},$
where $m^{\prime}=m\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-m^{1/2+\epsilon}$.
@@ 3.2 Full conformal and jackknife+
Let $\mu_{\beta}:\mathcal{X}\to\mathbb{R}$ denote a predictor function parameterized by $\beta\in\mathbb{R}^{p}$. By a slight abuse of notation, let the map $T:\cup_{n\geq 1}(\mathcal{X}\times\mathcal{Y})^{n}\to\mathbb{R}^{p}$ denote a training algorithm for estimating $\beta$, hence, $\hat{\beta}_{n}=T({\bf D}_{n})$. In this case, we have $\hat{\mu}_{{\bf D}_{n}}=\mu_{\hat{\beta}_{n}}$.
@@
$F_{Q}^{(n-1)}(t)$ $:=\operatorname{\mathbb{P}}_{Z_{1}\sim Q}\left(\left|Y_{1}-\mu_{\overline{\beta}_{-1}}(X_{1})\right|\leq t\right).$
@@
$\operatorname{\mathbb{P}}\Bigg{(}P_{e}^{\,\text{J+}}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log{4}/{\delta}}+2C\right)\sqrt{\frac{B}{n}}+L_{Q,n-1}\,A(n,p,\epsilon)\Bigg{)}\leq\epsilon+\delta,$
@@
Using the same arguments as in the proof of this theorem, one can get a coverage bound for the CV+ as well. Unlike (2) which is meaningful only if the number of samples in each fold $\ell$ is large, the bound we present in the following corollary is suitable for cases where $\ell/n\to 0$.
@@
$\operatorname{\mathbb{P}}\Bigg{(}$ $P_{e}^{\,\text{CV+}}({\bf D}_{n})>\alpha+\sqrt{\frac{\log(2/\delta)}{2n}}+2\,\ell\,L_{n-\ell}\,\kappa_{2}\,c_{n-\ell}\bigg{(}\frac{1}{\kappa_{1}}+\sqrt{\frac{n}{2\kappa_{1}^{2}}\log\frac{2p}{\epsilon}}\bigg{)}\Bigg{)}\leq\epsilon+\delta.$
@@
$\operatorname{\mathbb{P}}\left(P_{e}^{\text{full}}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log(4/\delta)}+2C\right)\sqrt{\frac{B}{n}}+L_{Q,n}\,E(n,p,\epsilon)\right)\leq\epsilon+\delta.$
@@
4 Fully conditional conformal prediction
In this section, we study conformal prediction conditional on both training dataset and test bin.
@@ 5 Discussion
with $X_{\text{test}}\perp\!\!\!\perp{\bf D}_{n+m}$ and ${\bf D}_{n+m}=((X_{1},Y_{1}),...,(X_{n+m},Y_{n+m}))$. Although weaker than uniform stability, these parameters are yet not well-understood in a practical sense. To elaborate on the difference between this approach and uniform stability, we evaluate their resulting training-conditional bounds for the ridge regression under exchangeability.
@@
$\operatorname{\mathbb{P}}\Bigg{(}P_{e}({\bf D}_{n})>\alpha+O\big{(}n^{-1/2}\big{(}\sqrt{{\log(1/\delta)}}+$ $\sqrt{p\log(2p/\epsilon)}\big{)}\big{)}\Bigg{)}\leq\epsilon+\delta.$